OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 15 degrees Celsius is. I have Part 1 done. I just need part 2. Please help.

OpenStudy (anonymous):

Plug C=15 into the equation. So: 15=5/9(F-32) F-32=27 F=27+32 F=59 degrees

OpenStudy (anonymous):

thank you. c: